# Reductions in local certification

### v1

@@ 8. Final remarks
Since the lower bounds obtained in this paper are all based on NP-hardness reductions, all the applications presented so far relate to coNP-hard problems. It might be tempting to conjecture that all such problems have polynomial local complexity. We now observe that this is not the case.
Indeed, consider the property $\mathcal{P}$ defined as the set of connected graphs $G$ for which there exists a vertex $u\in V(G)$ and an integer $m\geqslant 2$ such that $u$ has degree $3m$ and the neighbors $v_{1},\ldots,v_{3m}$ of $u$ can be partitioned into $m$ triples $T_{1},\ldots,T_{m}$ such that, for all $i\in\{1,\ldots,m\}$, $\Sigma_{v\in T_{i}}\deg(v)=\frac{1}{m}\Sigma_{j=1}^{3m}\deg(v_{j})$. The problem of determining whether a graph satisfies $\mathcal{P}$ is NP-complete : indeed, there is a simple polynomial reduction from 3-partition where the inputs are encoded in unary (recall that 3-partition is strongly NP-complete), which is the following: on an instance :=$\mathcal{I}:=\{a_{1},\ldots,a_{3m}\}$ of 3-partition (where $m\geqslant 2$), we create the graph $G_{\mathcal{I}}$ consisting of a central vertex $u$ that has $3m$ neighbors $v_{1},\ldots,v_{3m}$, and for every $i\in\{1,\ldots,m\}$, $v_{i}$ has $a_{i}-1$ additional neighbors of degree 1. For every $i\in\{1,\ldots,n\}$, we have $\deg(v_{i})=a_{i}$. The graph $G_{\mathcal{I}}$ satisfies $\mathcal{P}$ if and only if we can partition the set of neighbors of $u$ into triples of vertices that have the same sum of degrees, so if and only if $\mathcal{I}$ is a valid instance of 3-partition. However, there is a local reduction with certificates of size $O(\log n)$ for $\overline{\mathcal{P}}$, where the certificate of each vertex is its degree. Then, every vertex can simply check that its set of neighbors cannot be partitioned into such triples.
@@
Let us give examples of reductions that apply to problems that are not coNP-hard. For some graphs $H$, it is known that certifying that a graph $G$ does not contain $H$ as an induced subgraph (we will call this property $H$-freeness) requires certificates of size $\Omega(n)$. For instance, it is the case when $H$ is a clique of size at least 4 [4, Theorem 15] [^2] The model considered in [4] is a bit different from local certification but their proof can be easily adapted to work in this model., or when $H$ is a path of length at least 7 [1, Theorem 7] [^3] Authors in [1] actually prove the $\Omega(n)$ lower bound for $P_{7}$-graphs but their proof can be easily adapted to paths of length $\ell$ for every $\ell\geqslant 7$, by adding a pending path of length $\ell-7$ to the vertex $v_{0}$ in their construction.. However, for every fixed graph $H$, determining if an $n$-vertex graph $G$ contains $H$ as an induced subgraph can be done globally in time $O(n^{V(H)})$, so in polynomial time. Here, we present how, starting from a lower bound on the local complexity of $H$-freeness, we can prove a lower bound for $f(H)$-freeness, for some functions $f$ on graphs. For instance, let $k\geqslant 2$ and let $f$ be the following function: for every graph $G$, $f(G)$ is the graph obtained by replacing each vertex of $G$ by a clique of size $k$ and each edge of $G$ by a complete bipartite graph $K_{k,k}$ (this operation is sometimes called the lexicographic product of $G$ with $K_{k}$). It can be easily checked that, for every graphs $G$ and $H$, $G$ is $H$-free if and only if $f(G)$ is $f(H)$-free. Moreover, this transformation defines a local reduction of constant local expansion, where for each vertex $u\in V(G)$, the sets $V_{u}$ and $C_{u}$ are both equal to the clique of size $k$ replacing $u$ in $V(f(G))$. Thus, if $H$-freeness requires certificates of size $\Omega(n)$, $f(H)$-freeness requires certificates of size $\Omega(n)$ by Theorem 5.2. For instance, if $H$ is a path of length 7 and if $k=2$, we obtain an $\Omega(n)$ lower bound for the local complexity of $H^{\prime}$-freeness where $H^{\prime}$ is the following graph:
@@
If $f$ consists in adding a pending vertex to each vertex, it can easily be checked that $G$ is $H$-free if and only $f(G)$ is $f(H)$-free, and this defines again a local reduction of constant local expansion. So for instance, if $H$ is a path of length 7, we obtain an $\Omega(n)$ lower bound for the local complexity of $H^{\prime}$-freeness where $H^{\prime}$ is the following graph:
@@
If $f$ consists in adding a pending path of length $t$ to every vertex, if $i,j$ are two integers such that $i+j\geqslant 6$, it can easily be checked that $G$ contains no path of length $i+j+1$ if and only if $f(G)$ contains no claw with a central vertex and three branches of length $(t,t+i,t+j)$. This transformation defines again a local reduction of constant local expansion. For instance, with $t=3$, $i=2$, $j=4$, we obtain a $\Omega(n)$ lower bound on the local complexity of $H^{\prime}$-freeness where $H^{\prime}$ is the claw with branches of size $(3,5,7)$, that is, the following graph:

### v2

@@ 8. Final remarks
Since the polynomial lower bounds obtained in this paper are all based on NP-hardness reductions, all the applications presented so far relate to coNP-hard problems. It might be tempting to conjecture that all such problems have at least polynomial local complexity. We now observe that this is not the case, by giving a coNP-hard problem with logarithmic complexity.
Indeed, consider the property $\mathcal{P}$ defined as the set of connected graphs $G$ for which there exists a vertex $u\in V(G)$ and an integer $m\geqslant 2$ such that $u$ has degree $3m$ and the neighbors $v_{1},\ldots,v_{3m}$ of $u$ can be partitioned into $m$ triples $T_{1},\ldots,T_{m}$ such that, for all $i\in\{1,\ldots,m\}$, $\Sigma_{v\in T_{i}}\deg(v)=\frac{1}{m}\Sigma_{j=1}^{3m}\deg(v_{j})$. The problem of determining whether a graph satisfies $\mathcal{P}$ is NP-complete : indeed, there is a simple polynomial reduction from 3-partition where the inputs are encoded in unary (recall that 3-partition is strongly NP-complete), which is the following: on an instance :=$\mathcal{I}:=\{a_{1},\ldots,a_{3m}\}$ of 3-partition (where $m\geqslant 2$), we create the graph $G_{\mathcal{I}}$ consisting of a central vertex $u$ that has $3m$ neighbors $v_{1},\ldots,v_{3m}$, and for every $i\in\{1,\ldots,m\}$, $v_{i}$ has $a_{i}-1$ additional neighbors of degree 1. For every $i\in\{1,\ldots,n\}$, we have $\deg(v_{i})=a_{i}$. The graph $G_{\mathcal{I}}$ satisfies $\mathcal{P}$ if and only if we can partition the set of neighbors of $u$ into triples of vertices that have the same sum of degrees, so if and only if $\mathcal{I}$ is a valid instance of 3-partition. However, the local complexity of $\overline{\mathcal{P}}$ is $O(\log n)$: indeed, for each vertex $u$, the prover can just write the degree of $u$ in its certificate. Then, every vertex can simply check that the degree written in its certificate is correct and that its set of neighbors cannot be partitioned into such triples.
@@
Let us give examples of reductions that apply to problems that are not coNP-hard. For some graphs $H$, it is known that certifying that a graph $G$ does not contain $H$ as an induced subgraph (we will call this property $H$-freeness) requires certificates of size $\Omega(n)$. For instance, it is the case when $H$ is a clique of size at least 4 [5, Theorem 15] [^2] The model considered in [5] is a bit different from local certification but their proof can be easily adapted to work in this model., or when $H$ is a path of length at least 7 [1, Theorem 7] [^3] Authors in [1] actually prove the $\Omega(n)$ lower bound for $P_{7}$-free graphs but their proof can be easily adapted to paths of length $\ell$ for every $\ell\geqslant 7$, by adding a pending path of length $\ell-7$ to the vertex $v_{0}$ in their construction.. However, for every fixed graph $H$, determining if an $n$-vertex graph $G$ contains $H$ as an induced subgraph can be done globally in time $O(n^{|V(H)|})$, so in polynomial time. Here, we present how, starting from a lower bound on the local complexity of $H$-freeness, we can prove a lower bound for $f(H)$-freeness, for some functions $f$ on graphs. For instance, let $k\geqslant 2$ and let $f$ be the following function: for every graph $G$, $f(G)$ is the graph obtained by replacing each vertex of $G$ by a clique of size $k$ and each edge of $G$ by a complete bipartite graph $K_{k,k}$ (this operation is sometimes called the lexicographic product of $G$ with $K_{k}$). It can be easily checked that, for every graphs $G$ and $H$, $G$ is $H$-free if and only if $f(G)$ is $f(H)$-free. Moreover, this transformation defines a local reduction of constant local expansion, where for each vertex $u\in V(G)$, the sets $V_{u}$ and $C_{u}$ are both equal to the clique of size $k$ replacing $u$ in $V(f(G))$. Thus, if $H$-freeness requires certificates of size $\Omega(n)$, $f(H)$-freeness requires certificates of size $\Omega(n)$ by Theorem 5.2. For instance, if $H$ is a path of length 7 and if $k=2$, we obtain an $\Omega(n)$ lower bound for the local complexity of $H^{\prime}$-freeness where $H^{\prime}$ is the graph depicted in Figure 20.
Figure 20. The strong product of a path and a $K_{2}$.
@@
If $f$ consists in adding a pending vertex to each vertex, it can easily be checked that $G$ is $H$-free if and only $f(G)$ is $f(H)$-free, and this defines again a local reduction of constant local expansion. So for instance, if $H$ is a path of length 7, we obtain an $\Omega(n)$ lower bound for the local complexity of $H^{\prime}$-freeness where $H^{\prime}$ is the graph depicted in Figure 21.
Figure 21. A comb of length 7.
@@
If $f$ consists in adding a pending path of length $t$ to every vertex, if $i,j$ are two integers such that $i+j\geqslant 6$, it can easily be checked that $G$ contains no path of length $i+j+1$ if and only if $f(G)$ contains no claw with a central vertex and three branches of length $(t,t+i,t+j)$. This transformation defines again a local reduction of constant local expansion. For instance, with $t=3$, $i=2$, $j=4$, we obtain a $\Omega(n)$ lower bound on the local complexity of $H^{\prime}$-freeness where $H^{\prime}$ is the claw with branches of size $(3,5,7)$, depicted in Figure 22.
Figure 22. The claw with branches of size $(3,5,7)$.
Finally, let us mention some problems for which we were not able to apply our reduction framework. Let $G$ be an $n$-vertex graph, and let $S_{n}$ denote the set of all bijections $V(G)\to\{1,\ldots,n\}$. The bandwidth of $G$, denoted by $\beta(G)$, is defined as follows: $\beta(G):=\min_{\pi\in S_{n}}\max_{\{u,v\}\in E(G)}|\pi(u)-\pi(v)|$. The Bandwidth problem takes as input a graph $G$ and an integer $k$, and outputs whether $G$ has bandwidth at most $k$. We say that $G$ is a disk graph (resp. a unit-disk graph) if there exists a family of disks (resp. disks of diameter 1) in the plane such that $G$ is isomorphic to the intersection graph of these disks. The Disk Graph (resp. Unit-Disk Graph) problem takes as input a graph $G$, and outputs whether $G$ is a disk graph (resp. a unit-disk graph).
It is known that the Bandwidth, Disk Graph and Unit-Disk Graph problems are all NP-complete [20, 16, 3]. From the perspective of the framework introduced in this paper, it is natural to ask, for each of these problems, if there is a polynomial lower bound for the complementary property (where, for the Bandwidth problem, the integer $k$ is given as an input to the vertices). Unfortunately, we were not able to apply our local reduction framework and obtain polynomial lower bounds, either because existing NP-completeness reductions are not sufficiently local, or because they create too many vertices. Note also that, for each property to which we applied our reduction framework (summarized in Table 1), the complementary property, which is NP-complete, is also easy to certify (in the sense that certificates of constant or logarithmic size are sufficient). Surprisingly, here there is no straightforward logarithmic certification scheme for the Bandwidth, Disk Graph and Unit Disk Graph properties, which suggests that these properties may have a different behavior in the perspective of local certification. In fact, it is shown in [4] that the local complexity of Unit Disk Graph is $\Omega(n^{1-\delta})$ for any $\delta>0$. For the Bandwidth problem, a natural way to certify it would be to give to every vertex $u$ the integer $\pi(u)$ in its certificate. However, to certify that the function $\pi$ written in the certificates is injective, no upper bound better than $O(n)$ is known [2]. The previous observations thus raise the following problems.
Problem 8.1.
What is the local complexity of the complementary properties of Disk Graph and Unit-Disk Graph ?
Problem 8.2.
What is the local complexity of the Bandwidth property and of its complement ?
Acknowledgements
We thank the reviewers of the conference version of the paper for their helpful comments and suggestions.